# Solving the Poisson equation using coupled Markov chains

### v3

@@ B.2 Proof of Theorem 3
Lemma 29.
(Rio 1993, Theorem 1.1) Let $X$ and $Y$ be integrable random variables such that
@@
We follow the same strategy as in Douc et al. (2018, Lemma 21.4.3). By Lemma 30 with $v=P^{\ell}h$, we obtain
@@
for use in Lemma 29. It follows that
@@
By Lemma 31, it follows that
@@
Without loss of generality, assume $\pi(h)=0$. By Theorem 27, we have $g_{\star}\in L_{0}^{1}(\pi)$ since $\kappa>1$ and $m>2\kappa/(\kappa-1)>\kappa/(\kappa-1)$. Then, by Proposition 26 and Lemma 32 we obtain $\pi(h\cdot g_{\star})<\infty$. For the CLT, we appeal to Maxwell & Woodroofe (2000), for which it is sufficient to show that
@@
where $V_{n}f=\sum_{k=0}^{n-1}P^{k}f$, and $v(P,h)$ is then equal to $\lim_{n\to\infty}n^{-1}\mathbb{E}_{\pi}\left[\left\{\sum_{k=1}^{n}h(X_{k})\right\}^{2}\right]$. We find using Lemma 32,
@@
where we define $a_{k}=\min\{1,\mathbb{E}_{\pi\otimes\pi}[\tau^{\kappa}]^{\frac{m-2}{m}}(k+1)^{-\kappa\frac{m-2}{m}}\}$ by Proposition 26. Since $(a_{k})$ is non-increasing, we may deduce that
@@ B.3 Unbiased approximation of fishy functions
$\leq\left\|\sum_{i=0}^{\infty}\left|U_{i}\right|\mathbb{I}(N>i)+\left|V_{i}\right|\mathbb{I}(N>i)\right\|_{L^{p}}$
$\leq\sum_{i=0}^{\infty}\left\|\left|U_{i}\right|\mathbb{I}(N>i)\right\|_{L^{p}}+\left\|\left|V_{i}\right|\mathbb{I}(N>i)\right\|_{L^{p}}$
@@
The following lemma employs dominated convergence to justify the interchange of expectation of infinite sum and thereby ensure that $G_{y}(x)$ is indeed an unbiased estimator of $g_{y}(x)$.
Lemma 34.
@@
Since $m>\kappa/(\kappa-1)$, Theorem 27 provides that $g_{\star}\in L_{0}^{1}(\pi)$, and hence for $\pi$-almost all $x$ and $y$ we have
@@
The following lemma is used several times to ensure that expectations of functions of $X_{t}$ are uniformly bounded in $t$ under reasonable conditions. In the first case, the conclusion is a result of stability properties of $\pi$-invariant Markov chains at almost all points, while in the second case regularity is imposed by ensuring that $\mu$ cannot place too much probability in possibly problematic regions; for example it guarantees that $\phi\in L^{1}(\mu)$. It is possible that one can weaken the condition in the second statement to ${\rm d}\mu/{\rm d}\pi\in L^{p}(\pi)$ for some $p\geq 1$ but this would require stronger conditions on $\phi$ and hence complicate subsequent results.
Lemma 35.
@@
By assumption, $P$ is $\pi$-irreducible and from Proposition 26 it is aperiodic. Since $\pi(\phi)<\infty$, we may define $f=1+\phi\geq 1$ and $\pi(f)<\infty$. By the $f$-norm ergodic theorem (Meyn & Tweedie 2009, Theorem 14.0.1), for $\pi$-almost all $x$, $P^{t}f(x)$ is finite for all $t\geq 0$ and
@@
If $p\geq 1$ and $\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$, $\delta_{0}=p/\kappa$ and $\delta_{1}=p/m$ are in $(0,1)$ with $\delta_{0}+\delta_{1}<1$.We may therefore use Lemma 33 with $U_{t}=\left|h(X_{t})\right|$ and $V_{t}=\left|h(Y_{t})\right|$ to deduce that
@@
from which (B.7) follows. For the lack-of-bias property, $m>\kappa/(\kappa-1)$ implies $1>\frac{1}{m}+\frac{1}{\kappa}$, so the RHS of (B.7) is finite for $p=1$, and (B.6) holds for $\gamma$-almost all $(x,y)$. Since $\mathbb{E}_{\gamma}[\tau^{\kappa}]<\infty$ implies $\mathbb{P}_{x,y}(\tau<\infty)$ for $\gamma$-almost all $(x,y)$, we deduce by Lemma 34 that $\mathbb{E}\left[G_{y}(x)\right]=g_{y}(x)$ for $\gamma$-almost all $(x,y)$. It follows that
@@
Theorem 37.
Under Assumption 2, let $h\in L^{m}(\pi)$ for some $m>\kappa/(\kappa-1)$. Let $p\geq 1$ satisfy $\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$, and $\zeta$ and $\psi$ be as defined in Lemma 36. Let $\gamma$ be a probability measure with $\gamma_{1}=\gamma(\cdot\times\mathbb{X})$ and $\gamma_{2}=\gamma(\mathbb{X}\times\cdot)$.
@@
All of the parts are deduced from Lemma 36. For the first part, $\psi(h,m,\delta_{z})$ is finite for $\pi$-almost all $z$ by Lemma 35. For the second part, we add to this that $\mathbb{E}_{x,y}[\tau^{\kappa}]<\infty$ for $\pi\otimes\pi$-almost all $(x,y)$ by Assumption 2. For the third part, $\psi(h,m,\delta_{y})$ is finite for $\pi$-almost all $y$ and $\psi(h,m,\gamma_{1})^{m}\leq M\pi(\left|h\right|^{m})$ by Lemma 35, while
@@
for $\pi$-almost all $y$ by Assumption 2. For the fourth part, we add to this that $\psi(h,m,\gamma_{2})^{m}\leq M\pi(\left|h\right|^{m})$ by Lemma 35. ∎
@@ B.4 Unbiased approximation of $\pi(h)$
We next demonstrate that the approximation in Definition 15 is indeed unbiased and has finite $p$th moments under suitable conditions. The proof of Theorem 40 is an application of Theorem 37 and its statement can be compared with Middleton et al. (2020, Theorem 1), which treats the case $p=2$, for which we essentially arrive at the same condition for $\kappa$ and $m$. The lack-of-bias condition here is less demanding, and we deduce finiteness of higher moments of $H$ for $m$ sufficiently large.
In Propositions 4243 we establish that the properties obtained for $H$ may be deduced also for averages of lagged and offset estimators that are used in practice following Jacob et al. (2020). We use such unbiased estimators as part of the proposed asymptotic variance estimator UPAVE of Section 4.3. Finally, in Proposition 45 we show that subsampled estimators are also unbiased and have finite $p$th moments under the same conditions, because subsampling is an important aspect of UPAVE that controls its computational cost.
The following lemma guarantees that if an “independent initialization” is used, and ${\rm d}\mu/{\rm d}\pi\leq M$, then $\mathbb{E}_{\gamma}[\tau^{\kappa}]<\infty$ is guaranteed by Assumption 2, and similarly lack-of-bias results for $\gamma$-almost all $(x,y)$ may be deduced from lack-of-bias results for $\pi\otimes\pi$-almost all $(x,y)$.
Lemma 38.
Let $\gamma=\mu P\otimes\mu$ with $P$ a $\pi$-invariant Markov kernel. Then if ${\rm d}\mu/{\rm d}\pi\leq M$ then ${\rm d}\mu P/{\rm d}\pi\leq M$ and ${\rm d}\gamma/{\rm d}(\pi\otimes\pi)\leq M^{2}$.
@@
Lemma 39.
If $h,g_{\star}\in L^{1}(\pi)$ and $\mathbb{E}\left[G_{y}(x)\right]=g_{y}(x)$ for $\gamma$-almost all $(x,y)$, then
$\mathbb{E}\left[H\right]=\mathbb{E}\left[h(X_{0}^{\prime})+G_{Y_{0}^{\prime}}(X_{1}^{\prime})\right]=\pi(h).$
@@
$=\mu(h)-\mu(h-\pi(h))$
$=\pi(h).$
Theorem 40.
Under Assumption 2, let ${\rm d}\mu/{\rm d}\pi\leq M$, $\mathbb{E}_{\gamma}[\tau^{\kappa}]<\infty$, $h\in L^{m}(\pi)$ for some $m>\kappa/(\kappa-1)$. With $H$ defined in Definition 15, $\mathbb{E}\left[H\right]=\pi(h)$, and for $p\geq 1$ such that $\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$,
@@
Since $m>\kappa/(\kappa-1)$ and $\mathbb{E}_{\gamma}[\tau^{\kappa}]<\infty$, Theorem 37 gives $\mathbb{E}\left[G_{y}(x)\right]=g_{y}(x)$ for $\gamma$-almost all $(x,y)$ and Lemma 39 then implies that $\mathbb{E}[H]=\pi(h)$. By Minkowski’s inequality, Theorem 37, ${\rm d}\mu/{\rm d}\pi\leq M$, $m>p$, and $1\leq\zeta\left(\frac{(m-p)\kappa}{mp}\right)$ we obtain
$\mathbb{E}\left[\left|H\right|^{p}\right]^{\frac{1}{p}}$ $\leq\mathbb{E}\left[\left|h(X_{0})^{\prime}\right|^{p}\right]^{\frac{1}{p}}+\mathbb{E}\left[\left|G_{Y_{0}^{\prime}}(X_{1}^{\prime})\right|^{p}\right]^{\frac{1}{p}}$
@@
We now show how Theorem 40 can be used to extend the results to the approximations in Definition 16.
Remark 41.
The approximation may be viewed as the sum of $h(X_{k}^{\prime})$ and an unbiased approximation $G_{Y_{k}^{\prime}}^{(L)}(X_{k+L}^{\prime})$ of $g_{Y_{k}^{\prime}}^{(L)}(X_{k+L}^{\prime})=g_{\star}^{(L)}(X_{k+L}^{\prime})-g_{\star}^{(L)}(Y_{k}^{\prime})$, where $g_{\star}^{(L)}$ is the mean-zero solution of the Poisson equation for $(P^{L},h)$, i.e.
@@
The following shows that Theorem 40 holds for general $L\geq 1$ and $k\geq 0$, and in fact increasing either of these decreases the upper bound on the moments of the estimator.
Proposition 42.
@@
Recall that $a\vee b$ stands for the maximum and $a\wedge b$ for the minimum of $a$ and $b$. Since $\bar{P}^{L}$ is a coupling of $P^{L}$ with itself and $P^{L}$ is $\pi$-invariant, and following Remark 41, we seek to apply Theorem 40 to the approximation
$h(X_{k}^{\prime})+G_{Y_{k}^{\prime}}^{(L)}(X_{k+L}^{\prime}),$
@@
and we seek to verify that Assumption 2 holds for $\tau_{L,k}$ and obtain a finite bound for $\mathbb{E}_{\gamma_{k}^{(L)}}^{(L)}\left[\tau_{L,k}^{\kappa}\right]$, where $\mathbb{E}^{(L)}$ denotes expectation w.r.t. to the law of $(X^{(L,k)},Y^{(L,k)})$. In Theorem 40, we take $\mu=\pi_{0}P^{k}$.
@@
and therefore $\mathbb{E}_{\pi\otimes\pi}\left[\tau^{\kappa}\right]<\infty$ implies $\mathbb{E}_{\pi\otimes\pi}^{(L)}\left[\tau_{L,k}^{\kappa}\right]<\infty$. By Lemma 38, we have ${\rm d}\mu/{\rm d}\pi={\rm d}\pi_{0}P^{k}/{\rm d}\pi\leq M$ and ${\rm d}\gamma^{(L)}/{\rm d}(\pi\otimes\pi)\leq M^{2}$, and hence
@@
The following shows that the unbiased signed measure used in Section 4 and described in Appendix C is indeed unbiased for functions with suitably large moments when $\kappa$ is large enough, and that moments of averaged unbiased estimators are finite under the same conditions as for $H$.
Proposition 43.
@@
By Proposition 42, if $m>\kappa/(\kappa-1)>1$ then $\mathbb{E}[H_{t}]=\pi(h)$ for all $t\in\{k,\ldots,\ell\}$ and so $\mathbb{E}[H_{k:\ell}]=\pi(h)$. Using Minkowski’s inequality, we have
@@
from which we may conclude using the fact that the upper bound in (B.9) is non-increasing in $k$. ∎
We now demonstrate that estimators associated with subsampling the unbiased signed measure $\hat{\pi}$ described in Appendix C are unbiased and have finite $p$th moments under the same conditions as the standard estimator.
Lemma 44.
Let $\hat{\pi}=\sum_{i=1}^{N}\omega_{i}\delta_{Z_{i}}$ be the unbiased signed measure associated with $H_{k:\ell}$ in Proposition 43. Let the weights be as in the statement of Proposition 45 with $Z_{1},\ldots,Z_{N}$ corresponding to those points $X_{k}^{\prime},\ldots X_{\ell}^{\prime},X_{k+L}^{\prime},Y_{k}^{\prime},\ldots,X_{\tau-1}^{\prime},Y_{\tau-1-L}^{\prime}$ such that their weights are non-zero. In particular, there may be duplicate points in $Z_{1},\ldots,Z_{N}$. Then
@@
$=\frac{(\ell-k)/L+1}{\ell-k+1}$
$=\frac{1}{\ell-k+1}\left(1+\frac{\ell-k}{L}\right).$
Proposition 45.
Under Assumption 2, let $h\in L^{m}(\pi)$ for some $m>\kappa/(\kappa-1)$, ${\rm d}\pi_{0}/{\rm d}\pi\leq M$, $k,\ell\in\mathbb{N}$ with $k\leq\ell$, and $\hat{\pi}=\sum_{i=1}^{N}\omega_{i}\delta_{Z_{i}}$ be the unbiased signed measure associated with $H_{k:\ell}$ in Proposition 43:
$\frac{1}{\ell-k+1}\sum_{t=k}^{\ell}\delta_{X_{t}^{\prime}}+\sum_{t=k+L}^{\tau-1}\frac{\left\lfloor(t-k)/L\right\rfloor-\left\lceil\max(L,t-\ell)/L\right\rceil+1}{\ell-k+1}\left\{\delta_{X_{t}^{\prime}}-\delta_{Y_{t-L}^{\prime}}\right\},$
where $\tau=\inf\{t\geq L:X^{\prime}_{t}=Y^{\prime}_{t-L}\}$. Define for some $R\geq 1$,
$S_{R}=\frac{1}{R}\sum_{i=1}^{R}N\omega_{K_{i}}h(Z_{K_{i}}),$
where $K_{1},\ldots,K_{R}$ are conditionally independent ${\rm Categorical}\{\xi_{1},\ldots,\xi_{N}\}$ random variables satisfying
@@
for some constants $0<a\leq b<\infty$ independent of $N$. Then $\mathbb{E}[S_{R}]=\pi(h)$ and for $p\geq 1$ such that $\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$,
$\mathbb{E}\left[\left|S_{R}\right|^{p}\right]^{\frac{1}{p}}\leq a^{-1}b^{\frac{\kappa-p}{\kappa p}}\left\{\frac{1}{\ell-k+1}\left(1+\frac{\ell-k}{L}\right)\right\}\mathbb{E}\left[\left|N\right|^{\kappa}\right]^{\frac{m-p}{mp}}M^{\frac{1}{m}}\left\|h\right\|_{L^{m}(\pi)}\left\{\sum_{i=1}^{\infty}\frac{1}{i^{1+\varepsilon}}\right\}^{\frac{\kappa-p}{\kappa p}}<\infty,$
@@
The signed measure $\hat{\pi}$ is such that $\hat{\pi}(h)=H_{k:\ell}$. Hence, Proposition 43 may be applied to show that $\mathbb{E}[\hat{\pi}(h)]=\pi(h)$ when $m>\kappa/(\kappa-1)$. It follows that

### v4

@@ B.2 Proof of Theorem 3
Lemma B.6 (Rio 1993, Theorem 1.1).
Let $X$ and $Y$ be integrable random variables such that
@@
We follow the same strategy as in Douc et al. (2018, Lemma 21.4.3). By Lemma B.7 with $v=P^{\ell}h$, we obtain
@@
for use in Lemma B.6. It follows that
@@
By Lemma B.8, it follows that
@@
Without loss of generality, assume $\pi(h)=0$. By Theorem B.4, we have $g_{\star}\in L_{0}^{1}(\pi)$ since $\kappa>1$ and $m>2\kappa/(\kappa-1)>\kappa/(\kappa-1)$. Then, by Proposition B.3 and Lemma B.9 we obtain $\pi(h\cdot g_{\star})<\infty$. For the CLT, we appeal to Maxwell & Woodroofe (2000), for which it is sufficient to show that
@@
where $V_{n}f=\sum_{k=0}^{n-1}P^{k}f$, and $v(P,h)$ is then equal to $\lim_{n\to\infty}n^{-1}\mathbb{E}_{\pi}\left[\left\{\sum_{k=1}^{n}h(X_{k})\right\}^{2}\right]$. We find using Lemma B.9,
@@
where we define $a_{k}=\min\{1,\mathbb{E}_{\pi\otimes\pi}[\tau^{\kappa}]^{\frac{m-2}{m}}(k+1)^{-\kappa\frac{m-2}{m}}\}$ by Proposition B.3. Since $(a_{k})$ is non-increasing, we may deduce that
@@ B.3 Unbiased approximation of fishy functions
$\leq\left\|\sum_{i=0}^{\infty}\left|U_{i}\right|\mathds{1}(N>i)+\left|V_{i}\right|\mathds{1}(N>i)\right\|_{L^{p}}$
$\leq\sum_{i=0}^{\infty}\left\|\left|U_{i}\right|\mathds{1}(N>i)\right\|_{L^{p}}+\left\|\left|V_{i}\right|\mathds{1}(N>i)\right\|_{L^{p}}$
@@
The following lemma employs dominated convergence to justify the interchange of expectation and infinite sum, and thereby ensure that $G_{y}(x)$ is an unbiased estimator of $g_{y}(x)$.
Lemma B.11.
@@
Since $m>\kappa/(\kappa-1)$, Theorem B.4 provides that $g_{\star}\in L_{0}^{1}(\pi)$, and hence for $\pi$-almost all $x$ and $y$ we have
@@
The following lemma is used several times to ensure that expectations of functions of $X_{t}$ are uniformly bounded in $t$ under reasonable conditions. In the first case, the conclusion is a result of stability properties of $\pi$-invariant Markov chains started at almost all points, while in the second case regularity is imposed by ensuring that the initial distribution $\mu$ cannot place too much probability in possibly problematic regions; for example it guarantees that $\phi\in L^{1}(\mu)$. It is possible that one can weaken the condition in the second statement to ${\rm d}\mu/{\rm d}\pi\in L^{p}(\pi)$ for some $p\geq 1$, but this would require stronger conditions on $\phi$ and hence complicate subsequent results.
Lemma B.12.
@@
By assumption, $P$ is $\pi$-irreducible and from Proposition B.3 it is aperiodic. Since $\pi(\phi)<\infty$, we may define $f=1+\phi\geq 1$ and $\pi(f)<\infty$. By the $f$-norm ergodic theorem (Meyn & Tweedie 2009, Theorem 14.0.1), for $\pi$-almost all $x$, $P^{t}f(x)$ is finite for all $t\geq 0$ and
@@
If $p\geq 1$ and $\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$, $\delta_{0}=p/\kappa$ and $\delta_{1}=p/m$ are in $(0,1)$ with $\delta_{0}+\delta_{1}<1$. We may therefore use Lemma B.10 with $U_{t}=\left|h(X_{t})\right|$ and $V_{t}=\left|h(Y_{t})\right|$ to deduce that
@@
from which (B.7) follows. For the lack-of-bias property, $m>\kappa/(\kappa-1)$ implies $1>\frac{1}{m}+\frac{1}{\kappa}$, so the RHS of (B.7) is finite for $p=1$, and (B.6) holds for $\gamma$-almost all $(x,y)$. Since $\mathbb{E}_{\gamma}[\tau^{\kappa}]<\infty$ implies $\mathbb{P}_{x,y}(\tau<\infty)$ for $\gamma$-almost all $(x,y)$, we deduce by Lemma B.11 that $\mathbb{E}\left[G_{y}(x)\right]=g_{y}(x)$ for $\gamma$-almost all $(x,y)$. It follows that
@@
Theorem B.14.
Under Assumption 2, let $h\in L^{m}(\pi)$ for some $m>\kappa/(\kappa-1)$. Let $p\geq 1$ satisfy $\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$, and $\zeta$ and $\psi$ be as defined in Lemma B.13. Let $\gamma$ be a probability measure with $\gamma_{1}=\gamma(\cdot\times\mathbb{X})$ and $\gamma_{2}=\gamma(\mathbb{X}\times\cdot)$.
@@
All the parts are deduced from Lemma B.13. Recall that $\zeta$ is defined as in Lemma B.10. For the first part, $\psi(h,m,\delta_{z})$ is finite for $\pi$-almost all $z$ by Lemma B.12. For the second part, we add to this that $\mathbb{E}_{x,y}[\tau^{\kappa}]<\infty$ for $\pi\otimes\pi$-almost all $(x,y)$ by Assumption 2. For the third part, $\psi(h,m,\delta_{y})$ is finite for $\pi$-almost all $y$ and $\psi(h,m,\gamma_{1})^{m}\leq M\pi(\left|h\right|^{m})$ by Lemma B.12, while
@@
for $\pi$-almost all $y$ by Assumption 2. For the fourth part, we add to this that $\psi(h,m,\gamma_{2})^{m}\leq M\pi(\left|h\right|^{m})$ by Lemma B.12. ∎
@@ B.4 Unbiased approximation of $\pi(h)$
We next demonstrate that the approximation in Definition 15 is indeed unbiased and has finite $p$th moments under suitable conditions. The proof of Theorem B.17 is an application of Theorem B.14 and its statement can be compared with Middleton et al. (2020, Theorem 1), which treats the case $p=2$, for which we essentially arrive at the same condition for $\kappa$ and $m$. The lack-of-bias condition here is less demanding, and we deduce finiteness of higher moments of $H$ for $m$ sufficiently large.
In Propositions B.19B.20 we establish that the properties obtained for $H$ may be deduced also for averages of lagged and offset estimators, in Definition 16, that are used in practice. Finally, in Proposition B.22 we show that subsampled estimators are also unbiased and have finite $p$th moments under the same conditions. Subsampling is an important aspect of the proposed asymptotic variance estimator (UPAVE, Section 4.3). In this section we use $\mu$ as notation for a possible initial distribution of a chain evolving according to $P$. Later on this initial distribution will be set to either the user-specified distribution $\pi_{0}$ or to the distribution $\pi_{k}=\pi_{0}P^{k}$ for some integer $k$.
The following two lemmas guarantee that if an independent initialization from $\mu$ is used for coupled lagged chains, i.e. $\gamma=\mu P\otimes\mu$, and ${\rm d}\mu/{\rm d}\pi\leq M$, then $\mathbb{E}_{\gamma}[\tau^{\kappa}]<\infty$ is guaranteed by Assumption 2, and similarly lack-of-bias results for $\gamma$-almost all $(x,y)$ may be deduced from lack-of-bias results for $\pi\otimes\pi$-almost all $(x,y)$.
Lemma B.15.
Let $\mu$ be a probability measure on $(\mathbb{X},\mathcal{X})$ and $P$ a $\pi$-invariant Markov kernel. Then if ${\rm d}\mu/{\rm d}\pi\leq M$ then ${\rm d}\mu P/{\rm d}\pi\leq M$ and ${\rm d}(\mu P\otimes\mu)/{\rm d}(\pi\otimes\pi)\leq M^{2}$.
@@
Lemma B.16.
If $h,g_{\star}\in L^{1}(\pi)$ and $\gamma$ is a coupling of $\mu P$ and $\mu$ for which $\mathbb{E}\left[G_{y}(x)\right]=g_{y}(x)$ for $\gamma$-almost all $(x,y)$, then
$\mathbb{E}\left[H\right]=\mathbb{E}\left[h(X_{0}^{\prime})+G_{Y_{0}^{\prime}}(X_{1}^{\prime})\right]=\pi(h),$
where $H$ is in Definition 15 with $(X_{1}^{\prime},Y_{0}^{\prime})\sim\gamma$ and $X_{0}^{\prime}\sim\mu$.
@@
$=\mu(h)-\mu(h-\pi(h))=\pi(h).$
Theorem B.17.
Under Assumption 2, let $H$ and $\gamma$ be defined as in Definition 15, assume ${\rm d}\mu/{\rm d}\pi\leq M$, $\mathbb{E}_{\gamma}[\tau^{\kappa}]<\infty$, and $h\in L^{m}(\pi)$ for some $m>\kappa/(\kappa-1)$. Then, $\mathbb{E}\left[H\right]=\pi(h)$, and for $p\geq 1$ such that $\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$,
@@
Since $m>\kappa/(\kappa-1)$ and $\mathbb{E}_{\gamma}[\tau^{\kappa}]<\infty$, Theorem B.14 gives $\mathbb{E}\left[G_{y}(x)\right]=g_{y}(x)$ for $\gamma$-almost all $(x,y)$ and Lemma B.16 then implies that $\mathbb{E}[H]=\pi(h)$. By Minkowski’s inequality, Theorem B.14, ${\rm d}\mu/{\rm d}\pi\leq M$, $m>p$, and $1\leq\zeta\left(\frac{(m-p)\kappa}{mp}\right)$ we obtain
$\mathbb{E}\left[\left|H\right|^{p}\right]^{\frac{1}{p}}$ $\leq\mathbb{E}\left[\left|h(X_{0}^{\prime})\right|^{p}\right]^{\frac{1}{p}}+\mathbb{E}\left[\left|G_{Y_{0}^{\prime}}(X_{1}^{\prime})\right|^{p}\right]^{\frac{1}{p}}$
@@
We now show how Theorem B.17 can be used to extend the results to the approximations in Definition 16.
Remark B.18.
The approximation $H_{k}$ in Definition 16 may be viewed as the sum of $h(X_{k})$ and an unbiased approximation $G_{Y_{k}}^{(L)}(X_{k+L})$ of $g_{Y_{k}}^{(L)}(X_{k+L})=g_{\star}^{(L)}(X_{k+L})-g_{\star}^{(L)}(Y_{k})$, where $g_{\star}^{(L)}$ is the mean-zero solution of the Poisson equation for $(P^{L},h)$, i.e.
@@
The following shows that Theorem B.17 holds for general $L\geq 1$ and $k\geq 0$, and in fact increasing either of these decreases the upper bound on the moments of the estimator.
Proposition B.19.
@@
Recall that $a\vee b$ stands for the maximum and $a\wedge b$ for the minimum of $a$ and $b$. Since $\bar{P}^{L}$ is a coupling of $P^{L}$ with itself and $P^{L}$ is $\pi$-invariant, and following Remark B.18, we seek to apply Theorem B.17 to the approximation
$h(X_{k})+G_{Y_{k}}^{(L)}(X_{k+L}),$
@@
and we seek to verify that Assumption 2 holds for $\tau_{L,k}$ and obtain a finite bound for $\mathbb{E}_{\gamma_{k}^{(L)}}^{(L)}[\tau_{L,k}^{\kappa}]$, where $\mathbb{E}^{(L)}_{\gamma_{k}^{(L)}}$ denotes expectation w.r.t. to the law of $(X^{(L,k)},Y^{(L,k)})$. In Theorem B.17, we take $\mu=\pi_{0}P^{k}$.
@@
and therefore $\mathbb{E}_{\pi\otimes\pi}\left[\tau^{\kappa}\right]<\infty$ implies $\mathbb{E}_{\pi\otimes\pi}^{(L)}\left[\tau_{L,k}^{\kappa}\right]<\infty$. By Lemma B.15, we have ${\rm d}\mu/{\rm d}\pi={\rm d}\pi_{0}P^{k}/{\rm d}\pi\leq M$ and ${\rm d}\gamma^{(L)}/{\rm d}(\pi\otimes\pi)\leq M^{2}$, and hence
@@
The following shows that the unbiased signed measure (3.5) is indeed unbiased for functions with suitably large moments when $\kappa$ is large enough, and that moments of averaged unbiased estimators are finite under the same conditions as for $H$.
Proposition B.20.
@@
By Proposition B.19, if $m>\kappa/(\kappa-1)>1$ then $\mathbb{E}[H_{t}]=\pi(h)$ for all $t\in\{k,\ldots,\ell\}$ and so $\mathbb{E}[H_{k:\ell}]=\pi(h)$. Using Minkowski’s inequality, we have
@@
and we may conclude using the fact that the upper bound in (B.9) is non-increasing in $k$. ∎
Proof of Proposition 19.
We may write $H_{t}=h(X_{t})+B_{t}$ for a random variable $B_{t}$ defined on the right-hand side of (A.1) as $\sum_{j=1}^{\infty}h(X_{t+jL})-h(Y_{t+(j-1)L})$. Here $(X_{t+L},Y_{t})$ is a Markov chain with transition kernel $\bar{P}$, as generated by Algorithm 2.1. It follows that we may write
$\sqrt{\ell-k+1}\left(H_{k:\ell}-\pi(h)\right)=\frac{1}{\sqrt{\ell-k+1}}\sum_{t=k}^{\ell}h_{0}(X_{t})+\frac{1}{\sqrt{\ell-k+1}}\sum_{t=k}^{\ell}B_{t}.$
We now demonstrate that estimators associated with subsampling the unbiased signed measure $\hat{\pi}$ in (3.7) are unbiased and have finite $p$th moments under the same conditions as the standard estimator. We start with a technical lemma stating that the weights in $\hat{\pi}$ are bounded by expressions that depend on $k,\ell,L$ but not on $\tau$.
Lemma B.21.
Let $\hat{\pi}=\sum_{i=1}^{N}\omega_{i}\delta_{Z_{i}}$ be the unbiased signed measure in (3.7). Then
@@
$=\frac{(\ell-k)/L+1}{\ell-k+1}=\frac{1}{\ell-k+1}\left(1+\frac{\ell-k}{L}\right).$
Proposition B.22.
Under Assumption 2, let $h\in L^{m}(\pi)$ for some $m>\kappa/(\kappa-1)$, ${\rm d}\pi_{0}/{\rm d}\pi\leq M$, $k,\ell\in\mathbb{N}$ with $k\leq\ell$, and $\hat{\pi}=\sum_{i=1}^{N}\omega_{i}\delta_{Z_{i}}$ be the unbiased signed measure in (3.7). Define for some $R\geq 1$,
$S_{R}=\frac{1}{R}\sum_{i=1}^{R}\xi_{I_{i}}^{-1}\omega_{I_{i}}h(Z_{I_{i}}),$
where $I_{1},\ldots,I_{R}$ are conditionally independent ${\rm Categorical}\{\xi_{1},\ldots,\xi_{N}\}$ variables satisfying
@@
for some constants $0<a\leq b<\infty$ that may be functions of $k,\ell,L$ but not $\tau$. Then $\mathbb{E}[S_{R}]=\pi(h)$ and for $p\geq 1$ such that $\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$,
$\mathbb{E}\left[\left|S_{R}\right|^{p}\right]^{\frac{1}{p}}$$\leq$ $a^{-1}b^{\frac{\kappa-p}{\kappa p}}\left\{\frac{1}{\ell-k+1}\left(1+\frac{\ell-k}{L}\right)\right\}\mathbb{E}\left[\left|N\right|^{\kappa}\right]^{\frac{m-p}{mp}}$
×$\times M^{\frac{1}{m}}\left\|h\right\|_{L^{m}(\pi)}\left\{\sum_{i=1}^{\infty}\frac{1}{i^{1+\varepsilon}}\right\}^{\frac{\kappa-p}{\kappa p}}<\infty,$
@@
The signed measure $\hat{\pi}$ is such that $\hat{\pi}(h)=H_{k:\ell}$. Hence, Proposition B.20 may be applied to show that $\mathbb{E}[\hat{\pi}(h)]=\pi(h)$ when $m>\kappa/(\kappa-1)$. It follows that